OpenStudy (anonymous):

If f(x)=10^x show that f(x+h)-f(x) --------- = 10^x((10^h -1)/h) h

OpenStudy (across):

Let\[f(x)=10^x.\]Then\[\frac{f(x+h)-f(x)}{h}=\frac{10^x\cdot10^h-10^x}{h}=\frac{10^x(10^h-1)}{h}.\]

OpenStudy (across):

The key law here is\[a^{b+c}=a^b\cdot a^c.\]

OpenStudy (anonymous):

i still dont quite understand it tho...

OpenStudy (across):

It is very clear:\[f(x+h)=10^{x+h}=10^x\cdot10^h.\]